[Hallc_running] [New Logentry] RC Daily Update

pooser at jlab.org pooser at jlab.org
Sun Nov 18 10:15:01 EST 2018

Logentry Text:
RC report:
After the access to reset the shunt trip breakers at the beginning of day shift yesterday, beam was relatively stable and data collection was smooth for all shifts (https://logbooks.jlab.org/entry/3628153).  We began data collection for kinematic group 13 at the beginning of day shift yesterday and remained there until ~04:00 when the switch to kinematic group 14 was done preemptively as the lowest z-setting had not been taken.  

>From then on data collection for group 14 continued until 09:15 at which point we initiated a BCM calibration (https://logbooks.jlab.org/entry/3628163).  Once the BCM Calibration is completed we will continue with group 14.  Once group 14 is complete, we will have to switch the polarity of the SHMS and re-configure for the lowest z-setting in group 13 prior to moving onto group 16.

Unfortunately for Hall-A they are currently off as they are having issues with maintaining vacuum in their scattering chamber.  However, this is advantageous for us since we now have more available 15k supply so that we are not limited in cooling power on loop 2 as a result of the loop 3 JT valve being stuck (https://logbooks.jlab.org/entry/3627916).  Apparently this is a known hardware issue (as per Dave M.) and will require attention the next extended down.

Notices in effect for shift crews:
Shift leaders, please fill out the check list once per shift!
It is no longer required to fill out the standard.kinematics file online, it is now done automagically
Split the large desired charge between multiple runs (~1 hour per run)
Listed currents are estimates. For runs at less than 60uA, increase current until SHMS singles rates are between 400k-500k.
How-to power cycle PLC I/O chassis: https://logbooks.jlab.org/entry/3625862, https://logbooks.jlab.org/entry/3625900

Planned accelerator configuration changes:
Tuesday, Nov 20: CHL UPS battery change, should be noninvasive however, induce CHL anomaly is possible
Wednesday, Nov 21: Hall A goes offline and MCC wants to conduct high current beam study of energy spread at 40 uA.  It is unclear that we will be effected
Monday, Nov 26: Energy change to 1.86 GeV/pass, we will be moved over to 500 MHz laser.

Planned accesses to the hall:
Monday, Nov 26: aerogel change and HGC work needs to be coordinated with Walter K.

Special requests:
Vijay Kumar (Regina group) has requested a walkthrough during the next access
Jay B. would like to conduct another beam steering test (5 minute procedure) during the next polarity change.  He requests to only be contacted if the polarity change occurs during the hours of 0630 and 1930.
Carlos Munoz Camacho would like access to the hall to show Hall-C to two French engineers  working on the NPS project.  It should be a ~30 minute access, and is a low priority.  He can be contacted at 357-1065

There will be no RC meeting this weekend, the next meeting will take place at 16:30 on Monday 11/19.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3628165
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20181118/4040575a/attachment.html>

More information about the Hallc_running mailing list